OpenStudy (anonymous):

8/x + 15/x^2 = -1 A. x = 5, –3 B. x = –5, –3 C. x = –5, 3 D. x = 5, 3 Thanks.

OpenStudy (campbell_st):

multply each term by x^2 \[8x + 15 = x^2\] \[x^2 + 8x + 15 = 0\] factorise the quadratic (x+3)(x+5) = 0 now you can find your solutions

OpenStudy (anonymous):

I got D. Is that right?

OpenStudy (campbell_st):

nope... after factorising solve these equation x + 5 = 0 x + 3 = 0 they give the solutions

OpenStudy (anonymous):

B?

OpenStudy (campbell_st):

yes

OpenStudy (anonymous):

Okay. I always mess up with making the answer positive or negative. Thanks.